OpenStudy (anonymous):

Use guess-and-check to find two numbers that have the sum of 8 and a difference of 2.

OpenStudy (cydney_morgan):

5 and 3

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Calculate the volume of a sphere where the radius is 9 meters. 81 cubic meters 324 cubic meters 972 cubic meters 2916 cubic meters

OpenStudy (cydney_morgan):

V = 4/3*pi*r^3 V = 4/3* 3.14...* 729 meters = about 3058 cubic meters Rate

OpenStudy (anonymous):

whats the answer